OpenStudy (unicornsession):

Write the result in scientific notation (9 x 10 ^4) (8 x 10^6)

OpenStudy (anonymous):

\[9\times 8\times 10^{4+6}\] is a start, but not the finish

OpenStudy (unicornsession):

ok thank

OpenStudy (anonymous):

be careful

OpenStudy (anonymous):

the answer \[72\times 10^{10}\] is not right, because it is not in scientific notation

OpenStudy (unicornsession):

Got 7.2 x 10^11